Mr. and Mrs. Wang are both 40 years old. They have a pair of twin daughters and a son. The product of the ages of the children is 36. Find the possible ages of the children.
Answer by Earlsdon(6294)

Let's see what useful information we have!
The product of their ages is 36.
They have a pair of twin daughters.
Mr. & Mrs. Wang are 40 years old.
What are the possible ages of the children?
Let's first find the factors of 36 which will represent the possible ages of the three children:
1*1*36 = 36
1*2*18 = 36
1*3*12 = 36
1*4*9 = 36
1*6*6 = 36
2*2*9 = 36
3*3*4 = 36
Now select from this list the entries that have double (twin) integers.
1*1*36 = 36
1*6*6 = 36
2*2*9 = 36
3*3*4 = 36
The first entry doesn't qualify because if the parents are 40, they would not have a child who is 36.
This leaves the remaining entries:
1, 6, and 6.
2, 2, and 9.
3, 3, and 4.
as the possible ages.